# Automated System-level Testing of Unmanned Aerial Systems

### v1

@@ 5.2 Case Studies Description
For the first experiment, we use ArduCopter [^7] https://ardupilot.org/copter/ which is a part of the widely used open-source ArduPilot [^8] https://ardupilot.org/ardupilot/index.html project. The ArduCopter is used to operate a variety of multi-copter UAVs including a helicopter, tri-copter, quad-copter, hexacopter, etc. It supports a large variety of flight modes including manual, semi-autonomous, and fully autonomous. Its software is compatible with a wide range of hardware e.g., Pixhawk, CUAV v5 Plus, OpenPilot Revolution, etc. It has also been used for research purposes in a large number of experiments such as in [33][34], and [17].
@@
For the second experiment, we use an industrial case study representing the CDS of GCS (i.e., GCS-CDS). The GCS-CDS contains three important flight instruments as shown in Fig. 11. The first flight instrument represents an airspeed indicator (ASI). Fig. 11(a) shows an ASI displaying the airspeed value of $\approx$0 knots. The second flight instrument represents an altimeter. Fig. 11(b) shows an altimeter indicating the altitude value of $\approx$1660 feet above sea level (ASL). The third flight instrument represents a turn coordinator that is used to display the current turning direction and the bank angle of the UAV. Fig. 11(c) shows a turn coordinator displaying the turning direction towards the right and the bank angle of $\approx$3 degrees.
5.3 Experiments Setup
The modeling statistics of both case studies are shown in Table 1. For the first experiment using ArduCopter, we use the proposed UAS profile to model the domain and behavior of the ArduCopter. The ArduCopter domain model consists of six domain classes and 23 properties. The ArduCopter behavioral model comprises 12 flight states, 23 events, and 43 transitions. The domain model is shown in Fig. 6 and the flight behavioral model is shown in Fig. 7 and Fig. 8. The total number of applied stereotypes is 66. The expected behavior of the quadcopter is specified in the form of OCL constraints that are 48 in number (including general and state invariants). An excerpt from OCL constraints is presented in Listing 1. The domain model and OCL constraints are modeled by analyzing the source code of ArduCopter. We use the ArduPilot software-in-the-loop simulator that is built using the binaries of the same version of the ArduCopter source code.
For the second experiment using GCS-CDS, we use the UAS profile and CDS testing profile proposed in previous work [10]. The UAS profile is used to model the flight behavior of a UAV and the CDS testing profile to model the domain of the GCS-CDS. The GCS-CDS domain model consists of 41 classes and 216 properties. The ArduCopter behavioral model comprises 14 flight states, 25 events, and 45 transitions. The total number of applied stereotypes is 40. The expected behavior of GCS-CDS is specified in the form of OCL constraints that are 30 in number.
@@ 5.4 Experiments Execution
For the training phase, we use hyperparameters presented in Table 2. Since there are no standard guidelines for selecting hyperparameters for training, we select these hyperparameters based on the various experimental trials with different values of parameters and their combinations. We execute AITester and Random to analyze the training progress considering the randomness factor. Both AITester and Random approaches are executed on two machines with the same specifications, i.e., core i7 3.6 GHz processor, 64 GB RAM, 1 TB hard drive, NVIDIA GeForce®RTX 2080 Ti graphics card, and Windows 10 64-bit operating system. We set the number of episodes to 1000 for AITester and the same number of iterations for Random. At the end of the training session, a trained model for AITester is stored.
For the first experiment with ArduCopter, we used the trained AITester and ran it 100 times (Table 2). We also ran Random 100 times for the comparison. We use random as a baseline because the available approaches in literature either require manual test scenario creation, focus on testing a specific subsystem of UAV autopilot, or do not provide public tool support. Moreover, avionics testers commonly execute UAV flight scenarios with random actions and data using a flight simulator, a practice similar to random testing aimed at identifying system faults. Alternatively, if the testing focus is on uncovering security vulnerabilities, fuzz testing could serve as a comparative approach. While fuzz testing is security-focused [35], aiming to identify vulnerabilities through invalid or unexpected inputs, our evaluation does not aim for security testing of the UAS.
For the second experiment using GCS-CDS, we compare AITester with the baseline model-based CDS testing approach [10]. For the CDS testing approach, we repeated the experiment procedure mentioned in [10]. This experiment was executed on two machines with the same specifications used in the first experiment. We set the number of episodes to 100 (Table 2) for AITester and the same number of paths are used for the CDS testing approach. The overall execution time for each case study, across 100 episodes, was approximately 300 minutes. It is important to note that the execution time varies based on the machine specifications and the flight simulator being used. While we experimented with ArduPilot – a widely used open-source simulator – AITester is not tied to any specific simulator. It is adaptable and can seamlessly integrate with a faster simulator. To make our experimental results reproducible, we provide the toolset with the experiment setup and execution information at an online repository1.
@@ 5.5 Experiments Results and Analysis
Fig. 12 presents a comparison between AITester and Random based on the MAR for 1000 training episodes. The figure demonstrates MAR values (up to $\approx$650) that are computed over 1000 episodes using a larger window size [36], i.e., $n=350$. It can be observed that the MAR of AITester is high in most of the cases as compared to Random. MAR of AITester is greater than 50 in a majority of the episodes and in some cases, it is higher than 100, whereas Random maintains MAR $\approx$ 45 in most of the episodes. The MAR of AITester and Random is close in some cases. This happens when the AITester tries to explore (select random actions) the environment to create a balance between exploration and exploitation. From the training outcomes, it can be seen that AITester learns to perform the correct sequence of actions while violating the expected behavior of the SUT.
@@
To answer RQ1, we perform a comparison based on the expected behavior violations in the form of OCL constraints. AITester utilizes UAV flight state machines to explore different paths in a state machine. An important consideration is the exploration of diverse state machine paths during testing. To analyze the diverse exploration of UAV flight states, we use path diversity as a metric that is used in a similar context and different domains, e.g., for Web [37]. For RQ2, we analyze the statistically significant difference between the two approaches based on path diversity. We calculate path diversity using the method defined for state graphs [37]).
To analyze the results for RQ2 and perform a statistical comparison, we follow the guidelines defined by Arcuri and Briand [38]. We used the Wilcoxon signed-rank test with a significance level ($\alpha$) set at 0.05. The Wilcoxon signed-rank test is a non-parametric statistical test utilized for comparing two different sample sets. This test generates a p-value which is used to determine whether the difference between two samples is of statistical significance. A smaller p-value (normally $<=\alpha$) suggests that there is a statistically significant difference between the two samples. A larger p-value (normally $>\alpha$) indicates that there is no statistically significant difference between the two samples. To measure the magnitude of the difference, particularly to determine which sample is better, we utilized Cliff’s Delta [39]. Cliff’s Delta is a non-parametric effect size measure that quantifies the degree of difference between two samples. A value of 0 signifies that the two samples are identical. A positive value (up to +1) suggests that the first sample is better than the second sample. Conversely, a negative value (down to -1) indicates that the second sample is better than the first sample. The discussion on experiment results corresponding to each research question is given below.
@@
AltitudeHold
6964
4
4646
3
PositionHold
1780
4
1328
4
Climb
3820
4
2232
3
Decent
19140
4
2942
1
Landing
956
4
362
2
Loiter
872
4
374
2
Takeoff
696
4
690
4
@@
Statistical Tests AITester vs Random
AITester vs CDST Approach
Wilcoxon signed-rank 4.15298×$\times 10^{-07}$ ($<\alpha$)
2.93251×$\times 10^{-20}$ ($<\alpha$)
Cliff’s Delta Effect Size 1.0 $(Large)$
0.973144 $(Large)$
@@ 5.5.1 Results of Experiment on ArduCopter
To answer RQ1, the results of the comparison between AITester and Random testing based on violations of the expected behavior of the SUT (OCL violations) are presented in Table 3. The table shows the total and the unique number of OCL violations corresponding to different UAV flight states. One constraint may be violated multiple times when the UAV is in one state for a particular time. For example, if a UAV takes five seconds to reach the altitude specified for the Climb state and the environment changes are being observed after 500 milliseconds, one constraint for Climb state can violate 10 times. Therefore, repeated violation of the same OCL constraint is considered as one unique violation. To isolate unique OCL violations, we first grouped recurring violations by flight state, followed by a manual analysis to ascertain the uniqueness of each OCL violation. We keep track of all violations as well as the unique number of violations. From the table, it can be observed that the number of OCL violations done by the AITester in each flight state is greater than that of Random. It can be seen that the total number of OCL violations of AITester is approximately thrice as compared to Random. For some states such as Descent, Loiter, and Landing the AITester violates OCL constraints more than 50% as compared to Random. Moreover, for some states like Climb and AltitudeHold, the difference in the number of violations done by AITester is higher than that of Random. The difference between AITester and Random is noteworthy in the unique number of OCL violations. In some states (e.g., Climb and AltitudeHold,), the AITester violates slightly more OCL constraints as compared to Random. For the important flight states such as Descent, Loiter, and Landing, AITester violates all state invariants specified for these states, whereas Random can only violate a few. The total number of unique OCL violations done by AITester is more than 60% as compared to Random. The results suggest that AITester effectively violated several OCL constraints representing the expected behavior of ArduCopter. These violations of OCL constraints may lead to faults in ArduCopter. However, identifying the root causes of faults in such a system requires faults reproduction and faults localization [40] which is not in the scope of our approach.
RQ1 Result
AITester is effective in generating diverse test scenarios that lead to deviations from the expected behavior of ArduCopter.
@@
RQ2 Result
AITester is able to explore the diverse nature of flight paths from the behavioral model of UAV during testing.
@@
Armed
441
30
1
171
60
3
AltitudeHold
5744
1830
3
3107
3792
3
Approach
866
838
2
55
92
3
Circle
1902
2180
2
2521
3180
3
Climb
7599
5550
3
6051
6644
3
Cruise
5134
3620
3
3667
4078
3
Decent
6405
3914
3
17819
20103
3
Landing
963
960
3
62
61
3
Loiter
1837
907
2
283
426
3
Takeoff
788
405
2
63
11
3
@@ 5.5.2 Results of Experiment on GCS-CDS
To answer RQ1, the results show that the total number of OCL violations performed by the AITester is greater than that of the CDST approach. In most of the flight states, the AITester performed a higher number of OCL violations as compared to the CDST approach. The total number of unique OCL violations done by the AITester is 30 whereas the CDST approach can perform 24 total unique OCL violations. The analysis of flight state-wise unique OCL violations shows that the number of violations by AITester is greater than the CDST approach in some states, i.e., Armed, Approach, Circle, Loiter, and Takeoff.
The unique OCL violations performed by AITester lead to four different types of faults (Fig. 13) in GCS-CDS under test. These four types of faults are identified due to the violations of four OCL constraints shown in Listing 2. Due to the violation of OCL constraint C1, the first type of fault is found in ASI during Cruise flight state. According to the constraint C1, the airspeed value must be between 50 and 100. Fig. 13(a) shows that the airspeed value pointed by the ASI needle is 240 knots whereas the ASI tape shows a value of $\approx$44 knots. This indicates an inconsistency between the ASI needle and tape. Therefore, this type of fault is considered a CDS elements integration fault.
Listing 2: Failed OCL constraints corresponding to four faults in GCS-CDS
@@
The third and fourth types of faults are identified in Turn Coordinator during Circle and Loiter states due to the violation of OCL constraints C3 and C4 respectively. In the third type of fault, Turn Coordinator shows a bank angle out of the range specified by OCL constraint C3. Due to the out-of-range angle, the Turn Coordinator shows a high angle in Circle state as shown in Fig. 13(f). The fourth type of fault is identified due to the violation of OCL constraint C4. In this fault, while turning left during Decent (Fig. 13(g)) and Loiter (Fig. 13(h)) states, the Turn Coordinator shows values higher than -46. It was analyzed that the UAV was not flipped upside down during both states. This happened as the angle exceeded the maximum limit during Decent and Loiter states.
@@
RQ1 Result
AITester is effective in finding four potential faults in three different flight instruments of GCS-CDS.
For the RQ2, the results of the comparison between the AITester and CDST approach based on diversity are given in Table 4. The results of the Wilcoxon test suggest that p-value $<\alpha$, i.e., there is a statistically significant difference between the AITester and CDST approach. To identify which one is better, the results of the Cliff’s Delta effect size measure show a large effect size. The results of statistical tests suggest that the proposed approach (AITester) is better as compared to the CDST approach. That is, AITester outperforms the CDST approach in exploring the diverse nature of flight paths of UAV flight state machines.
RQ2 Result
AITester is capable of exploring diverse flight paths from the UAV flight behavioral model during the testing of GCS-CDS.

### v2

@@ 5.2 Case Studies Description
For the first experiment, we used ArduCopter [^7] https://ardupilot.org/copter/ which is a part of the widely used open-source ArduPilot [^8] https://ardupilot.org/ardupilot/index.html project. The ArduCopter supports operating a variety of multi-copter UAVs including helicopter, tri-copter, quad-copter, hexacopter, etc. It supports many flight modes including manual, semi-autonomous, and fully autonomous. Its software is compatible with a wide range of hardware e.g., Pixhawk, CUAV v5 Plus, OpenPilot Revolution, etc. It has also been used for research purposes in various experimental settings [36, 17].
@@
For the second experiment, we used an industrial case study representing the CDS of GCS (i.e., GCS-CDS). The GCS-CDS contains three important flight instruments as shown in Fig. 11. The first flight instrument represents an airspeed indicator (ASI). Fig. 11(a) shows an ASI displaying the airspeed value of $\approx$0 knots. The second flight instrument represents an altimeter. Fig. 11(b) shows an altimeter indicating the altitude value of $\approx$1660 feet above sea level (ASL). The third flight instrument represents a turn coordinator that is used to display the current turning direction and the bank angle of the UAV. Fig. 11(c) shows a turn coordinator displaying the turning direction towards the right and the bank angle of $\approx$3 degrees.
@@
5.3 Experiments Setup
The modeling statistics of both case studies are shown in Table 1. For the first experiment using ArduCopter, we used the proposed UAS profile to model the domain and behavior of the ArduCopter. The ArduCopter domain model consists of six domain classes and 23 properties. The ArduCopter behavioral model comprises 12 flight states, 23 events, and 43 transitions. The domain model is shown in Fig. 6 and the flight behavioral model is shown in Fig. 7 and Fig. 8. The total number of applied stereotypes is 66. The expected behavior of the quadcopter is specified in the form of OCL constraints that are 48 in number (including general and state invariants). We used the ArduPilot software-in-the-loop simulator which was built using the binaries of the same version of the ArduCopter source code.
For the second experiment using GCS-CDS, we used the UAS profile and CDS testing profile proposed in previous work [10]. The UAS profile was used to model the flight behavior of a UAV and the CDS testing profile was used to model the domain of the GCS-CDS. The GCS-CDS domain model consists of 41 classes and 216 properties. The ArduCopter behavioral model comprises 14 flight states, 25 events, and 45 transitions. The total number of applied stereotypes is 40. The expected behavior of GCS-CDS is specified in the form of OCL constraints that are 30 in number.
Listing 2: An excerpt of the OCL constraints used in our experiments
-- minimum battery level during Armed state
C1: context ArduCopter inv: self.oclIsInState(Armed) and self.battery.level>10
-- copter’s roll values range during PositionHold state
C2: context ArduCopter inv: self.oclIsInState(PositionHold) and self.attitude.roll>=0 and self.attitude.roll<=5
-- copter’s yaw and yaw_rate ranges during Circle state
C3: context ArduCopter inv: self.oclIsInState(Circle) and self.attitude.yaw>=-30 and self.attitude.yaw<=30
C4:context ArduCopter inv: self.oclIsInState(Circle) and self.attitude.yaw_rate>=0 and self.attitude.yaw_rate<=5
-- airspeed values range during Approach state
C5: context ArduCopter inv: self.oclIsInState(Approach) and self.airspeed>0 and self.airspeed<50
The constraints for our experiments were identified through a careful analysis of the Ardupilot documentation, as well as consultations with avionics testers to consider important testing scenarios. LABEL:lst:expconstraints presents a selection of OCL constraints utilized in our experiments. Constraint C1 defines the minimum battery level for ArduCopter in Armed state, aiming to ensure the UAV has sufficient battery to take off. If ArduCopter initiates takeoff with a low battery level and further depletes its battery during flight without sufficient power for a failsafe landing action, it could potentially result in a crash. Constraint C2 specifies roll angle range during PositionHold state to maintain the stability of the ArduCopter. If the roll angle surpasses the defined limit, it could destabilize the ArduCopter, potentially leading to collisions with obstacles. Constraints C3 and C4 define ranges for yaw angle and yaw rate during Circle state. Both yaw angle and rate are crucial in this state because violating these limits could potentially result in uncontrollable rotations, flipping over, loss of control, and possible collisions with obstacles. Constraint C5 specifies airspeed values range during Approach state. Exceeding the airspeed limit during this state may lead to a rapid and uncontrolled landing, or hit the ground with extreme force, potentially causing damage to both the copter and its surroundings. It is important to note that in our approach, these constraints do not force a UAV into unsafe scenarios. They mainly define the expected ranges for a specific flight state, and exceeding these ranges could signify deviations from the expected behaviors.
@@ 5.4 Experiments Execution
For the training phase, we used hyperparameters presented in Table 2. Since there are no standard guidelines for selecting hyperparameters for training, we selected these hyperparameters based on the various experimental trials with different values of parameters and their combinations. We executed AITester and Random to analyze the training progress considering the randomness factor. Both AITester and Random approaches were executed on two machines with the same specifications, i.e., core i7 3.6 GHz processor, 64 GB RAM, 1 TB hard drive, NVIDIA GeForce®RTX 2080 Ti graphics card, and Windows 10 64-bit operating system. We set the number of episodes to 1000 for AITester and the same number of iterations for Random. At the end of the training session, a trained model for AITester was stored.
For the first experiment with ArduCopter, we used the trained AITester and ran it 100 times (Table 2). We also ran Random 100 times for the comparison. We used random as a baseline because the available approaches in literature either require manual test scenario creation, focus on testing a specific subsystem of UAV autopilot, or do not provide public tool support. Moreover, avionics testers commonly execute UAV flight scenarios with random actions and data using a flight simulator, a practice similar to random testing aimed at identifying system faults. Alternatively, if the testing focus is on uncovering security vulnerabilities, fuzz testing could serve as a comparative approach. While fuzz testing is security-focused [37], aiming to identify vulnerabilities through invalid or unexpected inputs, our evaluation does not aim for security testing of the UAS.
For the second experiment using GCS-CDS, we compared AITester with the baseline model-based CDS testing approach [10]. For the CDS testing approach, we repeated the previous experiment procedure [10]. This experiment was executed on two machines with the same specifications used in the first experiment. We set the number of episodes to 100 (Table 2) for AITester and the same number of paths are used for the CDS testing approach. The overall execution time for each case study, across 100 episodes, was approximately 300 minutes. It is important to note that the execution time varies based on the machine specifications and the flight simulator being used. While we experimented with ArduPilot – a widely used open-source simulator – AITester is not tied to any specific simulator. It is adaptable and can seamlessly integrate with a faster simulator.
To make our experimental results reproducible, we provide the toolset with the experiment setup and execution information at an online repository1.
@@ 5.5 Experiments Results and Analysis
Fig. 12 presents a comparison between AITester and Random based on the MAR for 1000 training episodes. The figure demonstrates MAR values (up to $\approx$650) that are computed over 1000 episodes using a larger window size [38], i.e., $n=350$. It can be observed that the MAR of AITester is high in most of the cases compared to Random. MAR of AITester is greater than 50 in a majority of the episodes and in some cases, it is higher than 100, whereas Random maintains MAR $\approx$ 45 in most of the episodes. The MAR of AITester and Random is close in some cases. This happens when the AITester tries to explore (select random actions) the environment to create a balance between exploration and exploitation. From the training outcomes, it can be seen that AITester learns to perform the correct sequence of actions while violating the expected behavior of the SUT.
@@
To answer RQ1, we performed a comparison based on the deviations from the expected behavior, considering violations of OCL constraints. As AITester utilizes UAV flight state machines to explore different paths in a state machine, an important consideration is the exploration of diverse state machine paths during testing. To analyze the diverse exploration of UAV flight states, we used path diversity as a metric that was employed in a similar context and different domains, e.g., for Web [39]. For RQ2, we analyzed the statistically significant difference between the two approaches based on path diversity. We calculated path diversity using the method defined for state graphs [39]).
To analyze the results for RQ2 and perform a statistical comparison, we followed the guidelines defined by Arcuri and Briand [40]. We used the Wilcoxon signed-rank test with a significance level ($\alpha$) set at 0.05. The Wilcoxon signed-rank test is a non-parametric statistical test utilized for comparing two different sample sets. This test produces a p-value, which is used to determine if the difference between two samples holds statistical significance. A smaller p-value (normally $<=\alpha$) suggests that there is a statistically significant difference between the two samples. A larger p-value (normally $>\alpha$) indicates that there is no statistically significant difference between the two samples. To measure the magnitude of the difference, particularly to determine which sample is better, we utilized Cliff’s Delta [41]. Cliff’s Delta is a non-parametric effect size measure that quantifies the degree of difference between two samples. A value of 0 signifies that the two samples are identical. A positive value (up to +1) suggests that the first sample is better than the second sample. Conversely, a negative value (down to -1) indicates that the second sample is better than the first sample. The discussion on experiment results corresponding to each research question is given below.
@@
AltitudeHold 6964 4 4646 3
PositionHold 1780 4 1328 4
Climb 3820 4 2232 3
Descent 19140 4 2942 1
Landing 956 4 362 2
Loiter 872 4 374 2
Takeoff 696 4 690 4
@@
Statistical Tests AITester vs Random AITester vs CDST Approach
Wilcoxon signed-rank 4.15298×$\times 10^{-07}$ ($<\alpha$) 2.93251×$\times 10^{-20}$ ($<\alpha$)
Cliff’s Delta effect size 1.0 $(Large)$ 0.973144 $(Large)$
@@ 5.5.1 Results of Experiment on ArduCopter
To answer RQ1, the results of the comparison between AITester and Random testing based on violations of the expected behavior of the SUT (OCL violations) are presented in Table 3. The table shows the total and the unique number of OCL violations corresponding to different UAV flight states. One constraint may be violated multiple times when the UAV is in one state for a particular time. For example, if a UAV takes five seconds to reach the altitude specified for the Climb state and the environment changes are being observed after 500 milliseconds, one constraint for Climb state can violate 10 times. Therefore, repeated violation of the same OCL constraint is considered as one unique violation. To isolate unique OCL violations, we first grouped recurring violations by flight state, followed by a manual analysis to ascertain the uniqueness of each OCL violation. We keep track of all violations as well as the unique number of violations. From the table, it can be observed that the number of OCL violations done by the AITester in each flight state is greater than that of Random. It can be seen that the total number of OCL violations of AITester is approximately thrice compared to Random. For some states such as Descent, Loiter, and Landing the AITester violates OCL constraints more than 50% compared to Random. Moreover, for some states like Climb and AltitudeHold, the difference in the number of violations done by AITester is higher than that of Random. The difference between AITester and Random is noteworthy in the unique number of OCL violations. In some states (e.g., Climb and AltitudeHold,), the AITester violates slightly more OCL constraints compared to Random. For the important flight states such as Descent, Loiter, and Landing, AITester violates all state invariants specified for these states, whereas Random can only violate a few. The total number of unique OCL violations done by AITester is more than 60% compared to Random. The results suggest that AITester effectively violated several OCL constraints representing the expected behavior of ArduCopter. These violations of OCL constraints may lead to faults in ArduCopter. However, identifying the root causes of faults in such a system requires faults reproduction and faults localization [42] which is not in the scope of our approach.
RQ1 Result AITester is effective in generating diverse test scenarios that lead to deviations from the expected behavior of ArduCopter.
@@
RQ2 Result AITester is able to explore the diverse nature of flight paths from the behavioral model of UAV during testing.
@@
Armed 441 30 1 171 60 3
AltitudeHold 5744 1830 3 3107 3792 3
Approach 866 838 2 55 92 3
Circle 1902 2180 2 2521 3180 3
Climb 7599 5550 3 6051 6644 3
Cruise 5134 3620 3 3667 4078 3
Descent 6405 3914 3 17819 20103 3
Landing 963 960 3 62 61 3
Loiter 1837 907 2 283 426 3
Takeoff 788 405 2 63 11 3
@@ 5.5.2 Results of Experiment on GCS-CDS
To answer RQ1, the results show that the total number of OCL violations performed by the AITester is greater than that of the CDST approach. In most of the flight states, the AITester performed a higher number of OCL violations compared to the CDST approach. The total number of unique OCL violations done by the AITester is 30 whereas the CDST approach can perform 24 total unique OCL violations. The analysis of flight state-wise unique OCL violations shows that the number of violations by AITester is greater than the CDST approach in some states, i.e., Armed, Approach, Circle, Loiter, and Takeoff.
The unique OCL violations performed by AITester lead to four different types of faults (Fig. 13) in GCS-CDS under test. These four types of faults are identified due to the violations of four OCL constraints shown in LABEL:gcscds-violated. Due to the violation of OCL constraint C1, the first type of fault is found in ASI during Cruise flight state. According to the constraint C1, the airspeed value must be between 50 and 100. Fig. 13(a) shows that the airspeed value pointed by the ASI needle is 240 knots whereas the ASI tape shows a value of $\approx$44 knots. This indicates an inconsistency between the ASI needle and tape. Therefore, this type of fault is considered a CDS elements integration fault.
Listing 3: Failed OCL constraints corresponding to four faults in GCS-CDS
@@
The third and fourth types of faults are identified in Turn Coordinator during Circle and Loiter states due to the violation of OCL constraints C3 and C4 respectively. In the third type of fault, Turn Coordinator shows a bank angle out of the range specified by OCL constraint C3. Due to the out-of-range angle, the Turn Coordinator shows a high angle in Circle state as shown in Fig. 13(f). The fourth type of fault is identified due to the violation of OCL constraint C4. In this fault, while turning left during Descent (Fig. 13(g)) and Loiter (Fig. 13(h)) states, the Turn Coordinator shows values higher than -46. It was analyzed that the UAV was not flipped upside down during both states. This happened as the angle exceeded the maximum limit during Descent and Loiter states.
@@
RQ1 Result AITester is effective in finding four potential faults in three different flight instruments of GCS-CDS.
For the RQ2, the results of the comparison between the AITester and CDST approach based on diversity are given in Table 4. The results of the Wilcoxon test suggest that p-value $<\alpha$, i.e., there is a statistically significant difference between the AITester and CDST approach. To identify which one is better, the results of the Cliff’s Delta effect size measure show a large effect size. The results of statistical tests suggest that the proposed approach (AITester) is better compared to the CDST approach. That is, AITester outperforms the CDST approach in exploring the diverse nature of flight paths of UAV flight state machines.
RQ2 Result AITester is capable of exploring diverse flight paths from the UAV flight behavioral model during the testing of GCS-CDS.
Figure 14: A test path selected from experimental results, explored by AITester during the testing process. It illustrates the sequence of states, actions, and evaluations of OCL constraints at each step along the path. This path eventually led to locating a fault in GCS-CDS shown in Fig. 13(h) during the Loiter state due to the violation of an OCL constraint C4 (LABEL:gcscds-violated).
Figure 15: Transformation of the test path from Fig. 14 into an executable test script.
5.5.3 Test Paths in Experiments
During the experiments, AITester generated multiple test paths, with violations of OCL constraints occurring in various flight states. Fig. 14 illustrates one of the test paths selected from the experimental results. This test path corresponds to a fault observed in the Turn Coordinator of the GCS-CDS, as shown in Fig. 13(h). The fault occurred during the Loiter state due to the violation of an OCL constraint C4 (LABEL:gcscds-violated). While traversing this test path, AITester executed the Arm action when the UAV was in the Disarmed state, resulting in a transition to the Armed state. After the state change, AITester evaluated the constraints specific to the Armed state. Next, AITester performed Takeoff action which led to the TakeOff state, and proceeded with the evaluation of constraints associated with this state. In a similar manner, AITester selected the actions IncreaseAltitude, HoldAltitude, DecreaseAltitude, HoldAltitude, and Loiter. These actions guided the UAV through the Climb, AltitudeHold, Descent, AltitudeHold, and Loiter states, respectively. Simultaneously, AITester evaluated the constraints applicable to each state. One of the constraints for Loiter state was violated, indicating a deviation from the expected behavior. AITester continued the testing process and performed further actions to take the UAV into AltitudeHold, Descent, Descent, Landing, and finally to Disarmed states. Upon completing this process, the deviation reported by AITester was utilized to locate the fault.
Reproducing Test Paths. Each AITester-generated test path consists of a series of actions executed by AITester during the testing process. To reproduce the test paths generated by AITester, each action must be transformed into corresponding UAV flight operations. Fig. 15 shows the transformation of a test path, as depicted in Fig. 14, into an executable test script. Initially, for each action in the test path, a corresponding UAV flight operation call must be defined within the test script. For instance, Arm action would be transformed into arm() operation. Subsequently, for each UAV operation call, a script specific to the flight simulator must be defined. As illustrated in Fig. 15, the scripts for arm() and takeoff() operations correspond to the Ardupilot flight simulator. If a different flight simulator is used, scripts for each operation need to be specifically defined for that particular simulator. Ultimately, this procedure results in test scripts corresponding to the test paths. These scripts can be utilized to re-execute the test path. Note that the process of transforming test paths into test scripts can be automated using the CDST tookit [43].